Fermi National Laboratory

Accelerator Update for July 16, 2008 - July 18, 2008


Wednesday July 16

The day shift began with Operations monitoring the Tevatron (TeV) store 6300 (182.93E30) and the Recycler (RR) stash (88.0E10), with the Antiproton Source stacking (36.2E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0 using the colliding beams, and with the SciBooNE / MiniBooNE (BNB) and Meson MTest experiments taking beam.

At 8:33 AM, the Linac Klystron Vernier tripped off on a transformer water flow fault. It reset.

At 8:55 AM, a Linac RF station (LRF5) tripped off. It reset.

At 9:05 AM, the Booster Gradient Magnet Power Supply (GMPS) and BRF3 tripped off. They both reset.

Operations disabled the MTest beam at 10:25 AM and began preparing the Meson M01-M05 enclosures for access.

A MI expert began a study at 10:27 AM.

At 11 AM, the Booster West Anode Power Supply (WAPS) and BRF11 tripped off. They both reset.

At 11:13 AM, KRF3 tripped off. It reset.

At 12:20 PM, the Pbar Sweeper magnet “A” tripped off and would not reset. Operations contacted an expert who brought Sweeper magnet “B” online while he investigated.

A MI expert began a multibatch slipstacking study at 3:34 PM.

At 3:38 PM, a Pbar expert reported that one of the core 4-8 momentum system TWTs had failed and would not be repaired until Friday. The expert made adjustments to help with the core momentum cooling efficiency until the repair.

At 3:50 PM, a Linac Debuncher quadrupole magnet (D:QF) alarmed with a ground fault. Experts investigated. It reset.

Operations reported at 4:01 PM that the Recycler had stashed antiprotons five times during the day shift.

Operations established beam to NuMI at 4:36 PM.

Operations established beam to MTest at 6:02 PM.

At 9:08 PM, the TeV sector B2 wet engine failed. Operations terminated store 6300 to prevent a quench. Operators contacted cryo system experts at CHL to repair the wet engine.

Cryo system experts reported at 10:49 PM that the B2 wet engine’s motor had burned out.

At 11:29 PM, LRF3 tripped off. It reset.

Operations reported at midnight that the Recycler had stashed six times during the evening shift.

Thursday July 17

The midnight shift began with the TeV waiting on B2 wet engine repairs, with Operations monitoring the stash (408.6E10E10), with Pbar stacking (24.2E10), and with BNB and MTest taking beam.

Cryo system technicians reported that they had completed their repairs on the B2 wet engine at 1:09 AM.

At 1:35 AM, the NuMI beam permit tripped off on a target closed loop low-level expansion tank indication. Operations contacted experts.

Operations started up the TeV ramp at 1:57 AM.

Operations put the TeV through a dry squeeze at 2:13 AM.

Operations began preparing NuMI for an access at 2:17 AM.

Operations put the TeV into shot setup at 2:38 AM.

A Recycler expert began conducting drag rate measurements at 4:06 AM.

Operations established store 6302 at 4:16 AM with an initial luminosity of 288.33E30.

Operations turned off beam to BNB and began preparing them for an access at 5:28 AM.

Operations shut off all beam at 5:38 AM in preparation for the MI shutdown day.

Interlock experts began conducting BNB safety system tests at 5:42 AM.

Operations prepared Linac for an access at 5:54 AM.

Operations prepared the Meson M01-M05 enclosures for access at 5:57 AM.

Interlock experts began conducting Linac safety system tests at 6:05 AM.

A water expert reported at 6:58 AM that he had started the MI LCW heat exchanger work. (This work was the driving force for the shutdown.)

Interlock experts reported at 7:50 AM that they had completed testing the MI-12A & B safety systems.

Operations reported at 11:11 AM that the NuMI closed loop cooling problem was resolved.

Linac came out of access at 11:14 AM.

At 3:39 PM, Operations reported that KRF3 would not turn on. Linac experts investigated.

A power supply expert reported at 4:25 PM that the following jobs had been completed: the Buchholz relay for the MI-50 QD supply was replaced, the interphase transformer in E:Q847 was replaced, and the obsolete Recycler control cables were removed from the MI-20, 30, 40, and 50 enclosures.

A Booster expert reported at 4:53 PM that Booster came out of access, that BRF5 was operational, and that they were hipotting GMPS.

Linac experts reported at 8:03 PM that they were beginning a one-hour study on LRF5.

MI came out of access around 10:42 PM. Experts reported that MIRF7 was operational.

Friday July 18

The midnight shift began with Operations monitoring store 6302 (57.74E30) and the stack (112.3E10), with no stash, and with no experimental beam due to the MI shutdown.

At 12:13 AM, Operations reported that they were having problems with a Booster vacuum pump (B:IP31).

LCW experts reported at 1:28 AM that they had completed their work and pressure test at MI-10. The pressure test for MI-60 was in progress.

At 1:34 AM, a Booster vacuum pump (IP31) tripped off and caused BRF16 to trip off. BRF16 would not stay on; operators brought BRF19 online.

LCW experts reported at 2:42 AM that they were just finishing up with MI-60.

At 3:12 AM, KRF5 tripped off. It reset.

An LCW expert reported at 3:22 AM that he was beginning to turn on pond water pumps, starting at MI-10.

An LCW expert reported at 4:26 AM that all the pond water pumps were on except for the MI-20 #1 pump.

Operations established beam to Booster at 5:10 AM.

Operations turned on MI and established beam to Pbar at 5:53 AM.

Operations established beam to NuMI at 6:08 AM.

At 7:32 AM, an LCW expert reported that there was a problem with the Motor Control Center for MI-20. Operations contacted the Duty Electrician.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. However, a TeV sector C2 wet engine had a bearing failure and its repair will take about eight hours.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/18/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory